# Empirical Determination of Baseball Eras: Multivariate Changepoint Analysis in Major League Baseball

### v1

@@ 3.2.1 Attendance
We begin with attendance to be able to compare directly to the abundant literature using univariate change point methods on attendance data in professional sports. Attendance and other statistical measures’ change points can be found in Table 2 and Figure 2. We find similar change points to past work coinciding with the end of World War II (both for the mean in 1945, and the variance in 1944). Our analysis also identifies as change point in 1976, near the start of the Free Agency era and the expansion of the league to Toronto and (back to) Seattle. [17] identify an attendance change point in the mid-1960s followed by a very strong upward trend - particularly in the American League - beginning shortly thereafter. While they did not estimate a structural change during the 1960s time for the National League, there is evidence that an increasing attendance trend seems to have begun near our estimated 1976 change point. Figure 2 also shows this steepening attendance trend around this time. Finally, a variance change point was identified in 2008, which was the final year that baseball was played in both Yankee and Shea Stadium with both teams, the Yankees and the Mets, moving to stadiums with smaller capacities.
@@ 3.2.3 Home Runs and Walks
Figure 2: Change point of all statistics over the league for 120 years across 16 teams.
@@ 3.2.5 Strikeouts
3.3 Team Series: Dynasties and Collapses
@@
Figure 3: Change point location in time for each franchise throughout their history.

### v2

@@ 3.2.1 Attendance
We begin with attendance to be able to compare directly to the abundant literature using univariate change point methods on attendance data in professional sports. Attendance and other statistical measures’ change points can be found in Table 3 and Figure 3. We find similar change points to past work coinciding with the end of World War II (both for the mean in 1945, and the variance in 1944). Our analysis also identifies a change point in 1976, near the start of the Free Agency era and the expansion of the league to Toronto and (back to) Seattle. [17] identify an attendance change point in the mid-1960s followed by a very strong upward trend - particularly in the American League - beginning shortly thereafter. While they did not estimate a structural change during the 1960s time for the National League, there is evidence that an increasing attendance trend seems to have begun near our estimated 1976 change point. Figure 3 also shows this steepening attendance trend around this time. Finally, a variance change point was identified in 2008, which was the final year that baseball was played in both Yankee and Shea Stadium with both teams, the Yankees and the Mets, moving to stadiums with smaller capacities.
@@ 3.2.3 Home Runs and Walks
Figure 3: Change point of all metrics (Attendance, Walks, Home Runs, Runs, Stolen Bases, and Strikeouts) over the league for 120 years across 16 teams. The average value across all these teams is shown in bold with shaded standard deviation around it.
@@ 3.2.5 Strikeouts
3.3 Dynasties and Collapses: Modern Franchise
@@
Figure 4: Change point location in time for each franchise throughout their history, both in mean and variance if present. Gray bands represent the franchise’s lifespan into the modern era.